OpenStudy (anonymous):

what would be the simplified form of 1/b - 1/a over 1/b +1/a?

OpenStudy (anonymous):

b+1 over a^2(b-1)

OpenStudy (rajee_sam):

there are a lot of fractions right, so let us solve one by one. Let us just solve the numerator first. Your numerator is \[\frac{ 1 }{ b } - \frac{ 1 }{ a }\]

OpenStudy (anonymous):

b+a/b-a b-a/b+a a-b/a+b a+b/a-b

OpenStudy (rajee_sam):

@Mpinkham is wrong

OpenStudy (anonymous):

would it be b+a/b-a

OpenStudy (rajee_sam):

now there are 2 fractions with different Denominators. So we have to find LCM which is ab. So \[\frac{ (1 . a) - (1.b) }{ ab } = \frac{ a-b }{ ab}\]

OpenStudy (rajee_sam):

similarly if you simplify the denominator you wil get\[\frac{ a + b }{ ab }\]

OpenStudy (anonymous):

would it be a-b/a+b?

OpenStudy (rajee_sam):

\[\frac{ \frac{ a-b }{ ab } }{ \frac{ a+b }{ ab } } = \frac{ a-b }{ ab }\times \frac{ ab }{ a+b }\]

OpenStudy (rajee_sam):

Bingo yes

OpenStudy (anonymous):

thank you so much! Life saver, you are!